# Benchmarking Self-Supervised Methods for Accelerated MRI Reconstruction

### v2

@@ 2.1.6 Non-feedforward methods
Finally, we do not consider ground-truth-free methods that require lengthy iterative procedures or optimisation at inference-time, since these are impractical in real-world application, including Deep Image Prior methods (Darestani & Heckel, 2021), generative models e.g. AmbientGAN (Bora et al., 2018), diffusion methods (Daras et al., 2023; Kawar et al., 2024), implicit neural representation methods (Shen et al., 2024), or self-supervised plug-and-play training methods such as (Shafique et al., 2024).
@@ 5 Results
While SSDU qualitatively seems to remove some artifacts compared to no-learning, its metrics are very poor and drastically worse than no-learning; we explore why this is the case in Section 6.1 and show how to improve the results. Additionally, the inference-time modification inspired by Noise2Inverse from Eq. 5 does not seem to improve the results. Similarly, the weighted version weighted-SSDU from Eq. 7 performs badly too, and we show how to retrieve better results in Section 6.1.
@@ 6.1 Reproducing SSDU
We explore the seemingly poor performance of SSDU and attempt to reproduce the performance reported in (Yaman et al., 2020). First, we add 4 more unrolled iterations to the neural network from 3 to 7, increasing the training computational cost by more-than-double. Note that this means these results are no longer directly comparable with other methods reported in Section 5. We observe a considerable qualitative performance increase but not quantitatively in Figs. 4 and 3. Then, noting that SSDU is trained on further subsampled measurements ${\mathbf{M}}_{1}{\mathbf{y}}_{i}$ but tested on just the subsampled measurements ${\mathbf{y}}_{i}$, we normalise the model output to the range $[0,1]$. This improves the results, but removes any quantitative information in the pixel values. Finally, we standardise the output to the mean and standard deviation of the ground truth ${\mathbf{x}}$. While this now shows an improvement over no-learning, this requires ground-truth statistics and thus would no longer be fully self-supervised. We note that the Noise2Inverse inference time procedure also does not help, differing from comparisons observed in dynamic MRI (Wang & Davies, 2024a).
To explore the performance of weighted-SSDU, we do similar and increase the number of unrolled iterations. This massively boosts the performance of weighted-SSDU and it quantitatively outperforms the no-learning reconstruction, suggesting that although the theory holds agnostic of the neural network architecture, in practice, its performance highly depends on it.
@@
Weighted-SSDU (3) 6.94 .2970
Weighted-SSDU (7) 29.44 .7822
@@ Appendix E Additional results
SSDU 10.91 .4131
Noise2Inverse 11.22 .3969
Weighted-SSDU 8.00 .3614
VORTEX 27.97 .7023

### v3

@@ 2.1.6 Non-feedforward methods
Finally, we do not consider ground-truth-free methods that require lengthy iterative procedures or optimisation at inference-time, since these are impractical in real-world application, including Deep Image Prior methods (Darestani & Heckel, 2021), generative models e.g. AmbientGAN (Bora et al., 2018), diffusion methods (Daras et al., 2023; Kawar et al., 2024; Aali et al., 2024), implicit neural representation methods (Shen et al., 2024), or self-supervised plug-and-play training methods such as (Shafique et al., 2024).
@@ 5 Results
In this setup, the performance of SSDU is very poor and drastically worse than no-learning. However, we note that qualitatively it seems to remove some artifacts (Fig. 4), and we explore why this is the case in Section 6.1 and show how to improve the results. Additionally, the inference-time modification inspired by Noise2Inverse from Eq. 5 does not seem to improve the results. However, by adding the weighting from Eq. 7, weighted-SSDU performs much better; however, we notice that artifacts are still present in its reconstructions, reflected in its poorer SSIM.
@@ 6.1 Reproducing SSDU
We explore the seemingly poor performance of SSDU and attempt to reproduce the performance reported in (Yaman et al., 2020). First, we add 4 more unrolled iterations to the neural network from 3 to 7, increasing the training computational cost by more-than-double. Note that this means these results are no longer directly comparable with other methods reported in Section 5. We observe a considerable qualitative performance increase but not quantitatively in Figs. 4 and 3. Then, noting that SSDU is trained on further subsampled measurements ${\mathbf{M}}_{1}{\mathbf{y}}_{i}$ but tested on just the subsampled measurements ${\mathbf{y}}_{i}$, we normalise the model output to the range $[0,1]$. This improves the results, but removes any quantitative information in the pixel values. Finally, we standardise the output to the mean and standard deviation of the ground truth ${\mathbf{x}}$. While this now shows an improvement over no-learning, this requires ground-truth statistics and thus would no longer be fully self-supervised. We note that the Noise2Inverse inference time procedure also does not help, differing from comparisons observed in dynamic MRI (Wang & Davies, 2024a). We also observe a similar pattern in weighted-SSDU, whose performance improves as the number of unrolled iterations increases.
@@
Weighted-SSDU (3) 29.12 .8054
Weighted-SSDU (7) 31.64 .8381
@@ Appendix E Additional results
SSDU 10.91 .4131
Noise2Inverse 11.22 .3969
Weighted-SSDU 29.59 .7743
VORTEX 27.97 .7023